# Word problems and embedding-obstructions in cellular automata groups on groups

### v2

@@ 4.1 Dropping the centers
Suppose then $h^{\prime}(y)\in K^{\prime}\cdot y$, so $h^{\prime}(y)=\sigma_{\vec{v}}(y)$ for some $\vec{v}\in B_{rn}\cap K^{\prime}$. Since $h^{\prime}$ is not a shift by an element of $K^{\prime}$, we can find a pattern $P\in B^{B_{rn}}$ such that $h^{\prime}(z)\neq\sigma_{\vec{v}}(z)$ when $z\in[P]$. Positioning $z|B_{rn}$ and a central pattern of $y$ far from each other in the same $10rn\mathbb{Z}^{d}$-periodic point $x$, we obtain a point which $h^{\prime}$ does not shift by an element of $K^{\prime}$.
This concludes the proof, since as explained above, this gives a bound of the form $|H|\leq C^{n^{d}}!$ as needed in the theorem. ∎
@@ 5 Arboreous graph families
Definition 5.1.
Let $S$ be a finite set (think of it as the positive generating set of a free group). An $S$-labeled graph is a directed graph $G$ where each edge has a label from $S$ or $S^{-1}$. We say such a graph is good if
[…]
A graph is complete if for every vertex $x$, an outgoing edge exists with any label $s\in S^{\pm}$.
A weakly connected component of an $S$-labeled graph is just a connected component of the underlying undirected graph (with an undirected edge for each directed edge).
A good $S$-labeled graph can be thought of as a partial group action of the free group $F_{S}$. We also write $xs$ for the unique neighbor of $x$ where the edge has label $s$ (for $s\in S^{\pm}$), if this exists, and otherwise $xs$ is undefined.
@@
The proof is analogous to that of Theorem 3.1. We do not give it in detail, and only explain the main differences.
First, consider the case where $G$ contains a graph where for some vertex $h$, $h\{a,b,c\}^{*}$ is a copy of the ternary free monoid. In this case, we can directly use the arguments in the proof of Theorem 3.1, and perform dyadically accelerated funneling along the encoded ternary free monoids. For inputs $u$ where a witnessing tree exists (in the sense of Lemma A.3, proving $u\in L$), such a tree exists in some graph, and thus the $\mathrm{PAut}$-element is nontrivial.
If a witnessing tree does not exist, there is a slight subtlety, namely one may wonder if there can exist a graph where the tree is not injectively embedded, and this somehow leads to incorrectly thinking that it contains a witness. One can check that the logic of the proof does not assume that $hu\neq hv$ when $u,v\in\{a,b,c\}^{*}$ are two distinct words of the free monoid. The hard part is the inductive argument at the end, since there the same (first) track is acting both as control and as the permuted data. Here we note that the permutations are always commutatored with index checks, so the only situation where something nontrivial happens is that the index values are as they should along the computation tree. Thus, the controlling and permuted data is not mixed.
@@
Second, consider the general case, i.e. that we only have thickets with possibly polynomial paths between branchings in the trees. We still perform dyadically accelerated funneling, but now we need to think of the long non-branching paths as single edges, and find analogs of the automorphisms in the proof of Theorem 3.1 which can communicate over these paths.
The first thing we need is to have a way to check whether some node $h$ is branching, or more generally whether $hs=hs^{\prime}$ for $s,s^{\prime}\in F_{S}$. We can construct with the usual techniques first an automorphism that permutes the $\mathbb{Z}_{7}$-component at $h$ if the $\{0,1,2\}$-tracks at $hs$ and $hs^{\prime}$ have distinct values. Then we change the value at $hs$ and try again. One of these permutations succeeds if and only if $hs$ and $hs^{\prime}$ are distinct. The formulas from Section A.3 (or a variant of Equation 1) can now be used to obtain an automorphism that permutes the $\mathbb{Z}_{7}$-component at $h$ by a particular even permutation if and only if $hs\neq hs^{\prime}$.
Once we can check for equality of nodes, the proof is again similar to the proof of Theorem 3.1, but we check for an accepting tree on the tree obtained by contracting paths. We will construct versions of the $\phi$-maps that send information over paths where the tree is not branching, and goodness checks should similarly jump over paths. In the remainder of the proof, we explain how these variants can be constructed.
Note that the contracted paths are of length $n^{d}$ at most. We explain first how to deal with a single length $\ell$. A minor modification of Equation 1 gives an automorphism $f$ that performs $\pi$ in the $\mathbb{Z}_{7}$-component at $h$ if and only if $ha^{\ell}$ contains $s$. Namely, we simply replace the conjugating partial shifts of the third track by their $\ell$th power.
However, we need to check that the intermediate nodes on the path, i.e. $ha^{\ell^{\prime}}$ for $\ell^{\prime}<\ell$, do not branch, and that $ha^{\ell}$ does branch. This can be done since as we showed above, we have in our group automorphisms that permute $\mathbb{Z}_{7}$ in an arbitrary way depending on whether a given relative node is branching. Specifically, one can use commutator formulas and divide-and-conquer to obtain a polynomial-norm automorphism that cancels the effect of $f$ if the branching is not correct.
@@
To allow goodness checks to jump over paths, the argument is direct from Barrington’s theorem (taking the maps that check for branching nodes to be to be among the generators). ∎
6 PSPACE-hardness for general groups under the Gap Conjecture
@@
Let $b_{k}=|B_{k}|$ and suppose $G$ has at most $s$ generators. Suppose $x\in X$ has support contained in an $m$-ball, and has free orbit, assume $m$ is a window size for $X$. It is reasonably easy to show that there is a block map $\phi$ which maps $x$ to the characteristic function $\chi$ of $\{1_{G}\}$. Namely, consider the block map whose local rule identifies the central $B_{m}$-pattern of $x$ by writing $1$, and otherwise writes 00. If some shift of $gx$ also has this pattern in $B_{m}$, then the nonzero symbols of $x|g^{-1}B_{m}$ are in bijection with those of $x$ in $B_{m}$. Thus actually $g^{-1}x=x$ because these must contain all the nonzero symbols.
@@
The following is a variant of [29, Definition 1]; in that definition, the class of graphs was specialized to be a family of cyclic groups, but additionally a subshift was allowed on these groups. A subshift could be allowed on the graphs here as well, but the definition is already quite complicated, and we only need full shifts in our application.
@@
    To each $x\in X$ we can associate an $S$-labeled graph by taking nodes $V(X)=\{(g,j)\;|\;g\in G,0\leq j<m(x_{g})\}$, and for each $(g,j)\in V(x)$ with $v(x_{g})_{s,j}=(h,j^{\prime})$ adding an $s$-labeled edge from $(g,j)$ to $(gh,j^{\prime})$ (so we must have $(gh,j^{\prime})\in V(x)$). We require that this graph is good for all $x\in X$. Furthermore, the weakly connected components of these graphs are in $\mathcal{G}$.
@@
We say further that the simulation is complete if all the graphs are complete.
@@
for large enough $n$ (since $n^{2\beta d}>2$).
@@
Now the claim about tree branching linearly many times on each path amounts to showing that we can find $v\in\hat{L}$ with support size at least $n$. We show that indeed this is immediate from the cardinality of $\hat{L}$ (which is that of $|L|$), by some simple combinatorics.
Let $s$ be maximal number of non-$1$ coordinates in a word $v\in\hat{L}$. Then to describe words in $\hat{L}$ it suffices to list the $s$ positions of non-$1$ symbols (or a superset of them) and describe their contents. Thus we have
@@
To get an $(n,\{a,b,c\})$-thicket, we actually need the brachings to be ternary, i.e. along every path to a leaf, we should have many nodes with three children. This is not automatic, but by using paths of length at most $2$ of the tree (call these $2$-jumps) instead of the original edges (which correspond to length-$14k$ elements of $G$), it is possible to mimic the usual proof that $\{a,b,c\}^{*}$ embeds in $\{a,b\}^{*}$ as follows. First, as some preprocessing, step into a subtree so that every branching is binary, and drop every second branching so that there is a path of length at least $2$ between any consecutive branchings.
@@
By the choice of $k$, we can now position finite $k$-support patterns, namely central patterns of the set $F$ from Lemma 6.1, in the balls $B^{i}$ of the tree, where $|F|\geq C|B_{Ck}|^{C}$ for $C=\max(15,|B|)$. Distances in the tree of $B^{i}$s correspond to elements of $G$ with word norm at most $7k$, so double jumps correspond to elements with word norm at most $14k$, so it suffices to code four elements of $B_{14k}$ in each $B^{i}$, for which $|F|\geq C|B_{Ck}|^{C}$ is plenty since $C\geq 15\geq\max(4,14)$. We also still have room to store two symbols from the alphabet $B$, which is possible since $C\geq|B|^{2}$. Recall that a local rule can detect the positions of these patterns, and they do not overlap (by $3k$-separation), so automorphisms can freely change the $B$-symbol without messing up the encodings.
@@
To get not necessarily complete graphs, we could use $m(c)=1$ always. However, we want the graphs to be complete, while not removing any subgraphs, thus not removing the $3$-thickets whose existence we showed.
@@
Indeed the $3$-thickets still appear on the encoded graphs, on the first simulated nodes where edges go forward. ∎
Combining the above lemma with Theorem 5.1 gives the desired theorem.
@@
Let $G$ have at least stretched exponential growth, and let $X\subset A^{G}$ be any SFT such that some finite-support point has free orbit. Then $\mathrm{Aut}(X)$ has PSPACE-hard word problem.
@@
Let $G$ be a finitely-generated group with at least stretched exponential growth. Then $\mathrm{Aut}(A^{G})$ has PSPACE-hard word problem whenever $A\geq 2$.

### v3

@@ 4.1 Dropping the centers
Suppose then $h^{\prime}(y)\in K^{\prime}\cdot y$, so $h^{\prime}(y)=\sigma_{\vec{v}}(y)$ for some $\vec{v}\in B_{rn^{\prime}}\cap K^{\prime}$. Since $h^{\prime}$ is not a shift by an element of $K^{\prime}$, we can find a pattern $P\in B^{B_{rn^{\prime}}}$ such that $h^{\prime}(z)\neq\sigma_{\vec{v}}(z)$ when $z\in[P]$. Positioning $z|B_{rn^{\prime}}$ and a central pattern of $y$ far from each other in the same $10rn\mathbb{Z}^{d}$-periodic point $x$, we obtain a point which $h^{\prime}$ does not shift by an element of $K^{\prime}$. ∎
4.2 Embeddings into other groups
A version of Lemma 4.1 can be proved more generally for residually finite groups.
We recall the definition of systolic growth of a group [8].
Definition 4.9.
Let $G$ be a group. Its systolic growth with respect to a generating set $S$ is the function $f:\mathbb{N}\to\mathbb{N}$, where $f(n)$ is the minimal index of a subgroup $H$ such that the cosets $Hg$ are different for distinct $g\in B_{n}$.
Usually, one defines the systolic growth instead by taking $f^{\prime}(n)$ the minimal index of a subgroup $H$ such that $H\cap B_{n}=\{1_{G}\}$. However, it is easy to check that
$f^{\prime}(n)\leq f(n)\leq f^{\prime}(2n)$
so the functions are asymptotically equal.
Definition 4.10.
Let $G$ be a group and $X\subset A^{G}$ a subshift. Its periodization growth with respect to a generating set $S$ is the function $f:\mathbb{N}\to\mathbb{N}$ such that for any $p\in X|B_{n}$, there exists a point $x\in X$ with $x|B_{n}=p$ whose shift orbit has cardinality at most $f(n)$.
Lemma 4.11.
If $G$ has systolic growth $f$, then $A^{G}$ has periodization growth $f$.
Proof.
Let $p\in A^{B_{n}}$ be arbitrary. Let $H\leq G$ be a subgroup with index $f(n)$ such that the cosets $Hg$ for $g\in B_{n}$ are distinct. Extend $B_{n}$ to a set of right representatives $R$ for cosets of $H$, and pick any $q\in A^{R}$ with $q|B_{n}=p$. Define $x\in A^{G}$ by $x_{hk}=q_{k}$ for $h\in H$, $k\in R$. Then clearly $x|B_{n}=p$ and $x$ is $H$-periodic because $hx_{g}=x_{h^{-1}g}=x_{g}$ for $h\in H$. Therefore, $|Gx|\leq[G:H]=f(n)$. ∎
Lemma 4.12.
Let $G$ be a group, $X\subset A^{G}$ a subshift, and $\mathcal{G}=\langle S\rangle\leq\mathrm{Aut}(X)$, wheret $S$ is a finite generating set defining the word norm of $\mathcal{G}$. Suppose $X$ has periodization growth $f$. Then there exists $r\in\mathbb{N}$ such that whenever $\mathcal{H}\leq\mathcal{G}$ is a simple group containing a nontrivial element of norm at most $n$, we have $\mathcal{H}\leq\mathrm{Sym}(|A|^{f(rn)})$.
Proof.
Let $r$ be the maximal radius among generators in $S$. Suppose $h\in\mathcal{H}$ has norm $n$. Then the radius of $h$ is at most $rn$. Since $h$ is nontrivial, the local rule of $h$ maps some pattern $p\in X|B_{rn}$ to a symbol distinct from $p_{e}$.
Let $x\in X$ be a point with $x|B_{n}=p$ whose shift orbit has cardinality at most $f(n)$, and let $H$ be the stabilizer (of index at most $f(n)$) of $x$. Let
$P=\{y\in X\;|\;\forall h\in H:hy=y\}.$
Since the natural action of $\mathrm{Aut}(X)$ commutes with the shift action, in particular $\mathcal{H}$ stabilizes $P$ (as a set). We have $hx\neq x$ since $x|B_{n}=p$ and by the assumption on the local rule of $h$. Since $H$ is simple, its action on $P$ is faithful. We conclude that $H\leq\mathrm{Sym}(P)$, where of course $|P|\leq|A|^{f(rn)}$. ∎
Of course, this can be improved by taking into account the growth of the number of patterns in $X$ that appear in $H$-periodic points, where $H$ ranges over the group used to realize periodization growth.
We again state the previous lemma in terms of simplistic growth.
Lemma 4.13.
Let $G$ be a group, and let $X$ have periodization growth $f$. Then every finitely-generated subgroup of $\mathrm{Aut}(X)$ has simplistic growth at most $e^{f(n)}!$.
We show as an application that the group of $5$-dimensional cellular automata does not embed in the cellular automata group on the Heisenberg group.
Proposition 4.14.
Let $H_{3}$ be the Heisenberg group of upper triangular $3$-by-$3$ integer matrices with unit diagonal. then $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ does not embed in $\mathrm{Aut}(B^{H_{3}})$ for $D\geq 5$.
Proof.
The group $H_{3}$ has $n^{4}$ systolic growth [8]. Thus, $B^{H_{3}}$ has periodization growth $n^{4}$ by Lemma LABEL:lem:PeriodizationGrowth. By Lemma 4.5, $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ has a finitely-generated subgroup with simplistic growth at least $e^{n^{D}/\mathrm{poly}\log(n)}$ with $D\geq 5$. If $\mathrm{Aut}(A^{\mathbb{Z}^{D}})\leq\mathrm{Aut}(B^{H_{3}})$, then this finitely-generated subgroup would have simplistic growth at most $e^{n^{4}}!$ by Lemma 4.13, a contradiction since
$e^{n^{D}/\mathrm{poly}\log(n)}!\not\lesssim e^{n^{4}}!.$
More generally, every nilpotent group has polynomial systolic growth. In fact, in [9] it is shown that even the residual girth is polynomial, and this function bounds the systolic growth from above.
5 Groups with free submonoids
Groups with free submonoids admit explicit cones: binary trees along two generators. Furthermore, since the free submonoid on two generators contains the one on three generators, we have a free direction where we can store any additional information. We show that this implies a lower bound on simplistic growth. We also show that by a simple modification of the notion of successfulness, the same proof gives PSPACE-hardness of the word problem of the group of cellular automata. Both of these results are generalized later, after we interpret graphs in configurations more abstractly in Section 6.
5.1 Simplistic growth lower bound
Theorem 5.1.
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any alphabet. Then $\mathrm{Aut}(A^{G})$ has simplistic growth at least $e^{2^{\sqrt[\ell]{n}}}!$ for some $\ell$. Thus it has an f.g. subgroup that does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
Proof.
We consider the case $A=C\times\{a,b,?\}$ with $|C|\geq 6$. In the case of a general alphabet, the marker/conveyor belt argument is similar as in Lemma 4.4. Note also that in the interesting special case of a free group $G$, $\mathrm{Aut}(A^{G})$ actually embeds in $\mathrm{Aut}(B^{G})$ for nontrivial alphabets $A,B$ [2].
We give a lower bound on simplistic growth far surpassing that of any $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$. If there is a free monoid on two generators, there is also a free submonoid on three generators $a,b,c$. Given a configuration $x\in A^{G}$, we construct a graph for ripple catching. At $g\in G$ read the word $\{a,b,?\}^{n}$ at $\{g,gc,gc^{2},\ldots,gc^{n-1}\}$. If the word is of the form $u?^{k}$ where $u\in\{a,b\}^{n-k}$. Then we take $g$ as a node with color $u$. Note that this color is an analog of the index from the $\mathbb{Z}^{d}$ case, and allows a unique color for each position in a binary tree.
We have a $t$-labeled edge from $g$ to $gt$ if both nodes exist, for $t\in\{a,b\}$. We have a natural partial action of $\{a,b\}^{*}$ which appends the generator at the end of the vertex color, and we use the actions of $a,b$ as their successor relations. The corresponding rank function is $\rho(w)=|w|$. Note that a node $g\in G$ being successful means that its color is of the form $w$ for $w\in\{a,b\}^{k}$, and for each $t\in\{a,b\}$, the color of $gt$ is precisely $w?^{k-1}$ (nodes $k=0$ are not successful).
Again, $\mathrm{PAut}$ implements the generators of the ripple group for this graph (with respect to the rank function and partial action of $\{a,b\}^{*}$), so that the generators of the ripple group have word norm $\mathrm{poly}(n)$. The only difference is that instead of incrementing numbers (or more precisely, checking that some number is the successor of another), we now append a symbol at the end of a word. The latter is much easier to implement (indeed, it is a matter of comparing individual symbols).
Then by Lemma 3.6, there is an element of word norm $\mathrm{poly}(\mathrm{poly}(n))=\mathrm{poly}(n)$ which acts nontrivially exactly on full cones, which contain 00 everywhere except at the root. Again, cones are self-indexing in the sense that the vertex color uniquely identifies a root of the cone, so this is an element of the simple group of all permutations of $C$-configurations on self-indexing cones.
Now the cones have $12^{n+1}-$ 1 nodes, so the cardinality of this simple group is $|C|^{2^{n+1}}!/2$, so the simplistic growth $f$ of $\mathrm{Aut}(A^{G})$ satisfies $f(n)\gtrsim e^{2^{\sqrt[\ell]{n}}}!$, far surpassing the simplistic growth $e^{n^{d}}!$ in the $d$-dimensional case. ∎
5.2 PSPACE-hardness of the word problem
Theorem 5.2.
Proof.
We work exactly as above, but now in addition to knowing their coordinates in the cone, we include a configuration for a pair of Turing machines (the PSPACE-complete problem we reduce involves two Turing machines), and we include in the notion of successfulness of a node (in other words, in the definition of the successor relations of $a,b$ on the color) that the Turing machines are performing computations. (Note that the Turing machines are nondeterministic, so the successor relations do not come from a partial action of $\{a,b\}^{*}$ in this case.)
Ripple catching is give us an element that acts nontrivially exactly on the roots of full computation trees. Such an element is nontrivial if a full computation tree exists, which corresponds exactly to a positive instance of the problem we are reducing.
We explain this in more detail. If $G$ contains a free monoid on two generators, it contains a free monoid on three generators, say powers of $a,b,c\in G$ satisfy no nontrivial positive relations.
We will use alphabet $A=C\times B\times(S\cup(Q\times S))$ for our f.g. group, where $B=\{a,b,?\}$. Here, $Q$ with $|Q|=t$, and $S$ with $|S|=s$ are respectively the common sets of states and tape symbols of the nondeterministic Turing machines $T_{a},T_{b}$ from Lemma B.4. The lemma states roughly that the language of words $u_{\mathrm{init}}\in S^{n}$ such that there exists a binary tree of depth $n$ of configurations of the machines, where in one direction we take steps by $T_{a}$, and in the other by $T_{b}$, the root of the tree contains the initial configuration corresponding to $u_{\mathrm{init}}$, and both machines accept at all leaves.
Our reduction takes a word $u_{\mathrm{init}}$ of length $n=2^{k}$ (note that one should think of this as a polynomial size input whose size just happens to be of nice form, rather than an exponential length word), and constructs an automorphism that performs a permutation $\pi$ on the $C$-component of $A$ whenever it sees a tree of the form described in Lemma B.4 on the configuration, encoded in words read on the $B$-components, and with $q_{\mathrm{init}}u_{\mathrm{init}}\in QS^{n}$ as the initial configuration on the tree (where $q_{\mathrm{init}}$ is the common initial state of $T_{a},T_{b}$).
We construct this by ripple catching, so we need to describe a graph where full cones correspond to valid computations trees. At $g\in G$, read words $(u,w)$ in coordinates $(g,gc,\ldots,gc^{n-1})$ where $u\in B^{n}$, $w\in(S\cup(Q\times S))^{n}$ (ignoring the $C$-components). We include the node in the ripple catching graph if the following hold:
  • $u$ is as in the previous proof, i.e. all the ????-symbols are at the end,
  • $w$ is a valid Turing machine configuration (has exactly one head),
  • if $u=?^{n}$, then $w$ is the initial configuration corresponding to $u_{\mathrm{init}}$,
  • if $u\in\{a,b\}^{n}$, then $w$ is a final configuration.
We include an edge from $g$ to $ga$ with label $a$ if both are nodes.
We have $t\in\{a,b\}$ act as in the previous proof on the word $u$ (i.e. by appending itself), and in the Turing machine component $w$, $t\in\{a,b\}$ acts by taking a step of the machine $T_{t}$.
We now implement successfulness in $\mathrm{PAut}$ exactly as before. The only new component is that we need to check that the Turing machine steps are performed correctly, to determine whether a node is successful. This is possible by Lemma B.17. All in all, checking successfulness requires word length polynomial in $n$. Combined with ripple catching, we have an element of word norm $\mathrm{poly}(\mathrm{poly}(n))=\mathrm{poly}(n)$, which acts nontrivially if and only if a full cone exists, in other words if and only if $u_{\mathrm{init}}$ is in the language $L$. ∎
@@ 6 Arboreous graph families
Recall the definition of an $S$-labeled graph from Section 3. We need some definitions for $S$-labeled graphs. Such a graph is complete if all $S^{\pm}$-successors exist. A weakly connected component of an $S$-labeled graph is just a connected component of the underlying undirected graph (with an undirected edge for each directed edge), and an $S$-labeled graph is weakly connected if it has just one weakly connected component.
@@
The proof is analogous to that of Theorem 5.2. We use $A=C\times B$ with $|C|\geq 6$ and $B=\{a,b,?\}$.
First, consider the case where $G$ contains a graph where for some vertex $h$, $h\{a,b,c\}^{*}$ is a copy of the ternary free monoid. In this case, we can directly use the element constructed in the proof of Theorem 5.2, and perform ripple catching along the encoded free monoids $\{a,b\}^{*}$, interpreting indices in initial segments of the cosets $\langle c\rangle$. For inputs $u$ where a witnessing tree exists (in the sense of Lemma B.4, proving $u\in L$), such a tree exists in some graph, and thus the $\mathrm{PAut}$-element is nontrivial.
On the other hand, if a witnessing tree does not exist abstractly, one cannot be embedded in graphs of any shape. Thus, as long as the ripple group is correcly implemented, we will have a correct reduction. There is now a possible source of worry, namely we could certainly have $uc^{i}=vc^{j}$ for distinct $u,v\in\{a,b\}^{*}$ and $i,j$. Such reuse does not matter as long as we are only permuting $C$-symbols as a function of the $B$-track. For the implementation of $\phi_{\pi,c}$, the $B$-symbols need to be changed as a function of $C$. In this case, we are interested in permuting the $C$-component of the color at a node $h\in H$ (where $H\in\mathcal{G}$) if the color at some relative position $ht$ for $t\in\{a,b\}^{*}$ contains $c\in C$. The proof of Lemma LABEL:lem:PhiIsInPAut goes through verbatim for implementing this by partial shifts and symbol permutations.
@@
Second, consider the general case, i.e. that we only have thickets with possibly polynomial paths between branchings in the trees. We still perform ripple catching, but now along the graph where long non-branching paths are contracted to single edges, and we need to again show that the ripple group can be implemented in a finitely-generated subgroup.
The first thing we need is to have a way to check whether some node $h\in H$ is branching, or more generally whether $hs=hs^{\prime}$ for $s,s^{\prime}\in F_{S}$. We can construct with the usual commutator techniques first an automorphism that permutes the $C$-component at $h$ if the $B$-tracks at $hs$ and $hs^{\prime}$ have distinct values. Then we change the value at $hs$ and try again, and finally change the value back. One of these permutations succeeds if and only if $hs$ and $hs^{\prime}$ are distinct. The details are similar to Equation 1. The usual commutator formulas can then be used to obtain an automorphism that permutes the $C$-component at $h$ by a particular even permutation if and only if $hs\neq hs^{\prime}$.
Note that the contracted paths are of length $n^{d}$ at most. We explain first how to deal with a single length $\ell$. Another variant of Equation 1 gives an automorphism $f$ that performs $\pi$ in the $C$-component at $h$ if and only if $ha^{\ell}$ contains the symbol $c$. Namely, we simply replace the conjugating partial shifts of the third track by their $\ell$th power.
However, we need to check that the intermediate nodes on the path, i.e. $ha^{\ell^{\prime}}$ for $\ell^{\prime}<\ell$, do not branch, and that $ha^{\ell}$ does branch. This can be done since as we showed above, we have in our group automorphisms that permute $C$ in an arbitrary (even) way depending on whether a given relative node is branching. Specifically, one can use commutator formulas and divide-and-conquer to obtain a polynomial-norm automorphism that cancels the effect of $f$ if the branching is not correct.
@@
To allow successfulness checks to jump over paths, the argument is intuitively direct from Barrington’s theorem (taking the maps that check for branching nodes to be to be among the generators), in the sense that NC1 is certainly strong enough a computational model to allow skipping such polynomial length paths. Alternatively, one can work analogously as in the construction of the $\phi$-automorphisms, and for each $\ell$ and $0\leq j\leq n$ construct a polynomial word norm controlled permutation that checks a value in $ha^{\ell}c^{j}$, and with commutator formulas have it cancel out unless the non-branching path from $h$ has length exactly $\ell$, and finally compose this over all the polynomially many $\ell$. ∎
Theorem 6.5.
Let $\mathcal{G}$ be any arboreous path family. Then for some $A$, the group $\mathrm{PAut}_{\mathcal{G}}(A)$ has simplistic growth at least $e^{2^{\sqrt[\ell]{n}}}!$ for some $\ell$.
Proof.
The proof is essentially a direct combination of the above proof (for communication over long paths), and Theorem 5.1.
The simple group we implement can be described as follows: For every $(n,\{a,b,c\})$-thicket encoded in the configuration, in the self-indexing form from the lemma, we act on the bits on the branching positions. It is crucial here that the group does not act on all of the bits on the encoded thicket, as otherwise it could be a direct product of actions on thickets of different shapes (which can have different numbers of nodes), and not simple at all. The element of this group that we construct is the one that performs a particular even permutation at the root of such a thicket, assuming all the branching positions contain the bit 00.
This is immediately obtain from the above proof, by simply dropping the Turing machines and only checking the indices. ∎
7 Arboreous groups
@@
Let $b_{k}=|B_{k}|$ and suppose $G$ has at most $s$ generators. Suppose $x\in X$ has support contained in an $m$-ball, and has free orbit, assume $m$ is a window size for $X$, and assume $m\geq 10$. It is reasonably easy to show that there is a block map $\phi$ which maps $x$ to the characteristic function $\chi$ of $\{1_{G}\}$. Namely, consider the block map whose local rule identifies the central $B_{m}$-pattern of $x$ by writing $1$, and otherwise writes 00. If some shift of $gx$ also has this pattern in $B_{m}$, then the nonzero symbols of $x|g^{-1}B_{m}$ are in bijection with those of $x$ in $B_{m}$. Thus actually $g^{-1}x=x$ because these must contain all the nonzero symbols, thus $g=1_{G}$.
@@
The following is a variant of [28, Definition 1]; in that definition, the class of graphs was specialized to be a family of cyclic groups, but additionally a subshift was allowed on these groups. A subshift could be allowed on the graphs here as well, but the definition is already quite complicated, and we only need full shifts in our application.
@@
    To each $x\in X$ we can associate an $S$-labeled graph by taking nodes $V(X)=\{(g,j)\;|\;g\in G,0\leq j<m(x_{g})\}$, and for each $(g,j)\in V(x)$ with $v(x_{g})_{s,j}=(h,j^{\prime})$ adding an $s$-labeled edge from $(g,j)$ to $(gh,j^{\prime})$ (so we must have $(gh,j^{\prime})\in V(x)$). We require that this graph is good for all $x\in X$.
@@
    The weakly connected components of the graphs are in $\mathcal{G}$, and furthermore, any $H\in\mathcal{G}$ appears as a weakly connected graphs.
  4. 4.
@@
for large enough $n$ (since $22\beta d>$ 2).
@@
Now the claim about the tree branching linearly many times on each path amounts to showing that we can find $v\in\hat{L}$ with support size at least $n$. We show that indeed this is immediate from the cardinality of $\hat{L}$ (which is that of $|L|$), by some simple combinatorics.
Let $s$ be the maximal number of non-$1$ coordinates in a word $v\in\hat{L}$. Then to describe words in $\hat{L}$ it suffices to list the $s$ positions of non-$1$ symbols (or a superset of them) and describe their contents. Thus we have
@@
To get an $(n,\{a,b,c\})$-thicket, we actually need the brachings to be ternary, i.e. along every path to a leaf, we should have many nodes with three children. This is not automatic, but by using paths of length at most $2$ of the tree (call these $2$-jumps) instead of the original edges, it is possible to mimic the usual proof that $\{a,b,c\}^{*}$ embeds in $\{a,b\}^{*}$ as follows. First, as some preprocessing, step into a subtree so that every branching is binary, and drop every second branching so that there is a path of length at least $2$ between any consecutive branchings.
@@
By the choice of $k$, we can now position finite $k$-support patterns, namely central patterns of the set $F$ from Lemma 7.2, in the balls $B^{i}$ of the tree, where $|F|\geq C|B_{Ck}|^{C}$ for $C=\max(|B|^{2},15)$. Distances in the tree of $B^{i}$s correspond to elements of $G$ with word norm at most $7k$, so double jumps correspond to elements with word norm at most $14k$, so it suffices to code four elements of $B_{14k}$ in each $B^{i}$, for which $|F|\geq C|B_{Ck}|^{C}\geq|B_{14k}|^{4}$ is plenty. We also still have room to store two symbols from the alphabet $B$, which is possible since $C\geq|B|^{2}$. Recall that a local rule can detect the positions of these patterns, and they do not overlap (by $3k$-separation), so automorphisms can freely change the $B$-symbol without messing up the encodings.
@@
To get not necessarily complete graphs, we could use $m(c)=1$ always. However, we want the graphs to be complete, while not removing any subgraphs, thus not removing the $(n,\{a,b,c\})$-thickets whose existence we showed.
@@
Indeed the $(n,\{a,b,c\})$-thickets still appear on the encoded graphs, on the first simulated nodes where edges go forward. ∎
Combining the above lemma with Theorem 6.4 gives:
@@
Let $G$ have at least stretched exponential growth, and let $X\subset A^{G}$ be any SFT such that some finite-support point has free orbit. Then $\mathrm{Aut}(X)$ has a finitely-generated subgroup with PSPACE-hard word problem.
@@
Let $G$ be a finitely-generated group with at least stretched exponential growth. Then $\mathrm{Aut}(A^{G})$ has a finitely-generated subgroup with PSPACE-hard word problem whenever $A\geq 2$.
Combining the above lemma with Theorem 6.5 gives:
Theorem 7.7.
Let $G$ have at least stretched exponential growth, and let $X\subset A^{G}$ be any SFT such that some finite-support point has free orbit. Then $\mathrm{Aut}(X)$ has simplistic growth at least $e^{2^{\sqrt[\ell]{n}}}!$ for some $\ell$. In particular, $\mathrm{Aut}(X)$ has a finitely-generated subgroup that does not embed in $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any $d$.
Proof.
The lower bound on simplistic growth is from Theorem 6.5, and then we observe that $e^{2^{\sqrt[\ell]{n}}}!$ exceeds the simplistic growth upper bound $e^{n^{d}}!$ for $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$. ∎
Corollary 7.8.
Let $G$ be a finitely-generated group with at least stretched exponential growth. Then $\mathrm{Aut}(A^{G})$ has simplistic growth at least $e^{2^{\sqrt[\ell]{n}}}!$ for some $\ell$, whenever $A\geq 2$. In particular, it has a finitely-generated subgroup that does not embed in $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any $d$.
We can similarly obtain a lower bound on simplistic growth.